Shown below are six statements about functions. Match each statement to one of the functions shown below which BEST matches that statement.

1. limx→6+f(x) and limx→6−f(x) both exist and are finite, but they are not equal.
2. The graph of y=f(x) has vertical tangent line at (6,f(6))
3. limx→6−f(x)=−∞.
4. limx→6+f(x) exists but limx→6−f(x) does not.
5. limx→6f(x)=∞.
6. limx→6f(x) exists but f is not continuous at 6.

A. f(x)=⎧⎩⎨⎪⎪⎪⎪cos(1x−6)02x+24if x<6if x=6if x>6
B. f(x)=⎧⎩⎨⎪⎪2x024−2xif x<6if x=6if x>6
C. f(x)=⎧⎩⎨⎪⎪2x02x−24if x<6if x=6if x>6
D. f(x)=x−6−−−−−√3
E. f(x)=1(x−6)2
F. f(x)=1x−6
1
Expert's answer
2020-11-30T19:07:59-0500

1. "\lim\limits_{x\to 6^+}f(x)" and "\lim\limits_{x\to 6^-}f(x)" both exist and are finite, but they are not equal.

C. "f(x)=x = \begin{cases}
 2x &\text{if } x<6 \\
 0 &\text{if } x=6 \\
 2x-24 &\text{if } x>6
\end{cases}"



2. The graph of "y=f(x)" has vertical tangent line at "(6, f(6))"

D. "f(x)=\sqrt[3]{x-6}"



3."\lim\limits_{x\to 6^-}f(x)=-\infin"

F. "f(x)=\dfrac{1}{x-6}"



4."\lim\limits_{x\to 6^+}f(x)" exists but "\lim\limits_{x\to 6^-}f(x)" does not

A. "f(x)=x = \begin{cases}
 \cos(\dfrac{1}{x-6}) &\text{if } x<6 \\
 0 &\text{if } x=6 \\
 2x+24 &\text{if } x>6
\end{cases}"



5."\lim\limits_{x\to 6}f(x)=\infin"

E. "f(x)=\dfrac{1}{(x-6)^2}"



6. "\lim\limits_{x\to 6}f(x)" exists but "f" is not continuous at "6"

B. "f(x)=x = \begin{cases}
 2x &\text{if } x<6 \\
 0 &\text{if } x=6 \\
 24-2x &\text{if } x>6
\end{cases}"